# LANTERN: Accelerating Visual Autoregressive Models with Relaxed Speculative Decoding

### v1

@@ Abstract
Auto-Regressive (AR) models have recently gained prominence in image generation, often matching or even surpassing the performance of diffusion models. However, one major limitation of AR models is their sequential nature, which processes tokens one at a time, slowing down generation compared to models like GANs or diffusion-based methods that operate more efficiently. While speculative decoding has proven effective for accelerating LLMs by generating multiple tokens in a single forward, its application in visual AR models remains largely unexplored. In this work, we identify a challenge in this setting, which we term token selection ambiguity, wherein visual AR models frequently assign uniformly low probabilities to tokens, hampering the performance of speculative decoding. To overcome this challenge, we propose a relaxed acceptance condition referred to as LANTERN that leverages the interchangeability of tokens in latent space. This relaxation restores the effectiveness of speculative decoding in visual AR models by enabling more flexible use of candidate tokens that would otherwise be prematurely rejected. Furthermore, by incorporating a total variation distance bound, we ensure that these speed gains are achieved without significantly compromising image quality or semantic coherence. Experimental results demonstrate the efficacy of our method in providing a substantial speed-up over speculative decoding. In specific, compared to a naïve application of the state-of-the-art speculative decoding, LANTERN increases speed-ups by $\mathbf{1.75}\times$ and $\mathbf{1.76}\times$, as compared to greedy decoding and random sampling, respectively, when applied to LlamaGen, a contemporary visual AR model.
@@ 1 Introduction
Auto-Regressive (AR) models have recently gained significant traction in image generation (Ramesh et al., 2021; Chen et al., 2020; Tian et al., 2024; Sun et al., 2024) due to their competitive performance, often matching or even surpassing diffusion models (Ho et al., 2020; Rombach et al., 2022). Notable examples include iGPT (Chen et al., 2020), DALL-E (Ramesh et al., 2021), VAR (Tian et al., 2024), and LlamaGen (Sun et al., 2024), which showcase the potential of AR models in image generation. Moreover, recent studies like Team (2024); Lu et al. (2023) have demonstrated that AR modeling can handle multi-modal data, including language and images, within a single unified framework. Given the remarkable success of AR models in language modeling, leading to the era of large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023), it is anticipated that AR modeling will emerge as a dominant paradigm for unifying multiple modalities into a single model in the near future.
@@
Figure 1: Images generated by vanilla decoding (top) and lossy speculative decoding with our relaxed acceptance condition (with mean accepted length 2.40) (bottom) on the text-conditioned LlamaGen Stage II (Sun et al., 2024).
@@
  • Our experiments using LlamaGen (Sun et al., 2024) as the target and EAGLE-2 (Li et al., 2024a) as the drafter demonstrate significant speed-ups, improving from 1.28×1.28$\times$1.28 × to $\mathbf{2.25}\times$ in greedy decoding and from 0.93×0.93$\times$0.93 × to $\mathbf{1.64}\times$ in random sampling, compared to the naive application of EAGLE-2, without substantial performance drop in terms of image quality.
@@
2 Token Selection Ambiguity in Visual Autoregressive Models
In this section, we introduce a novel problem, termed token selection ambiguity, which imposes a major challenge for applying speculative decoding into visual AR models. In Section 2.1, we provide empirical results on naïve application of existing speculative decoding methods into visual AR models. Section 2.2 presents the token selection ambiguity problem and how this problem causes the failure of speculative decoding in visual AR models.
2.1 Naïve Speculative Decoding Fails in Visual AR Models
Table 1: Mean accepted length of naïve application of existing speculative decoding methods on the greedy decoding regime. LlamaGen-L is used as a drafter for Speculative Decoding (Leviathan et al., 2023). Numbers in parentheses present the mean accepted length on MT-Bench (Zheng et al., 2023) with Vicuna-7B (Zheng et al., 2023). Results on Vicuna-7B are taken from EAGLE-2 (Li et al., 2024a).
Methods Mean Accepted Length
LlamaGen-3B Vicuna-7B
(Sun et al., 2024) (Zheng et al., 2023)
Speculative Decoding (Leviathan et al., 2023) 1.40 (-40.6%) 2.36
Medusa (Cai et al., 2024) 1.38 (-45.2%) 2.52
EAGLE-2 (Li et al., 2024a) 2.11 (-57.6%) 4.98
While speculative decoding has been successful in LLMs (Leviathan et al., 2023; Cai et al., 2024; Li et al., 2024a), we observe that its naïve application to visual AR models fails to provide noticeable speed-up. As shown in Table 1, the naïve application results in limited efficacy compared to its performance in LLMs, with a significant reduction in the mean accepted length, dropping between 41% to 58% across methods. To better understand this limitation, we conduct a comparative analysis of the next-token probability distributions in both LLMs and vision AR models, using Vicuna-7B (Zheng et al., 2023) on the MT-Bench dataset (Zheng et al., 2023) and LlamaGen models (Sun et al., 2024) on MS-COCO captions (Lin et al., 2014b), respectively.
Our analysis reveals that the drafter in the visual AR models often struggles to accurately predict the target model’s output, leading to frequent misalignment between token predictions from the drafter and target model. Specifically, Figure 2(a) demonstrates that the drafter model fails to achieve high accuracy in predicting the target model’s output even in top-3 predictions, whereas its LLM counterpart shows considerably higher accuracy. Consequently, Figure 2(b) clearly illustrates the significant misalignment in token prediction distributions between the drafter and target model in visual AR models.
2.2 Token Selection Ambiguity Problem
(a) Drafter Test Accuracy
(b) Target - Drafter Alignment
@@ 2.2 Token Selection Ambiguity Problem
Figure 2: (a) Top-1 and top-3 accuracy of learned drafter model on Vicuna-13B, LlamaGen-XL (text-conditioned, stage I), and LlamaGen-3B (class-conditioned) for predicting the target model’s outputs. (b) Heat map of the drafter and target model’s probability distribution during inference on text-conditioned LlamaGen-XL Stage I model. For simplicity, we only visualize the union of the top-20 probable tokens for each model. (c) An average top-1 and top-10 probabilities on Vicuna-7B and LlamaGen-XL (text-conditioned, stage I)
To understand the reason for the failure of speculative decoding in visual AR models, we identify a unique problem that we term as the token selection ambiguity. This problem stems from the fundamental differences between language and image data, as well as the distinct tokenization strategies employed in their respective AR models.
In language models, tokens represent discrete units such as words or subwords, which follow structured and predictable sequences governed by grammar and syntax (Zipf, 1935). This results in a concentrated next token probability distribution, where the model assigns high confidence to the most likely token. In contrast, visual AR models process pixels or patches as tokens, leading to spatially continuous and highly complex sequences. Consequently, these models often exhibit broad and dispersed next-token probability distributions, reflecting greater uncertainty and ambiguity in predicting the next token. We define this phenomenon as the token selection ambiguity problem.
To empirically validate this, we compare the next token probability distributions of Vicuna-7B, a language model, and text-conditioned LlamaGen, a visual AR model. Using 80 responses from Vicuna-7B on the MT-Bench dataset and 1,000 images from LlamaGen based on MS-COCO validation captions, we observe a stark contrast in how the models assign probabilities. As shown in Figure 2(c), LlamaGen has significantly lower top-1 and top-10 probabilities on average compared to Vicuna-7B. This highlights the greater ambiguity in selecting the next token in visual AR models, where the model struggles to confidently prioritize one token over others.
The implication of this problem is critical for speculative decoding since speculative decoding relies on concentrated next-token distributions which enables a simpler drafter to predict and align with the target model. Therefore, the dispersed token distributions in visual AR models make it difficult for the drafter to accurately approximate the target model’s token probabilities. This misalignment leads to a lower acceptance rate of drafter tokens during speculative decoding, ultimately reducing its effectiveness.
@@ 3.1 Latent Proximity Permits Token Interchangeability
This property arises from the tokenization process unique to visual AR models. Unlike text tokenization, which is straightforward due to its discrete nature, images are spatially continuous, making tokenization more complex (Esser et al., 2021). To handle this, models like vector-quantized variational auto-encoders (VQVAE) (Van Den Oord et al., 2017) and vector-quantized generative adversarial networks (VQGAN) (Esser et al., 2021) are used to discretize the latent embeddings of images. These embeddings maintain a continuous mapping between changes in latent space and the visual semantics of the generated images (Kingma & Welling, 2022; Goodfellow et al., 2014; Karras et al., 2019). As a result, small shifts in latent space lead to minor shifts in the image, supporting the idea that tokens close in the latent space are effectively interchangeable.
Figure 3: Image generated by text-conditioned LlamaGen Stage II model (Sun et al., 2024) with five captions from MS-COCO (Lin et al., 2014b) validation set. The images are generated by either standard sampling method (top) or sampling with random replacement within 100-closest tokens in the latent space (bottom).
To demonstrate this concept empirically, we perform an experiment in which, after each token is sampled, it is re-sampled uniformly from the 100 closest tokens in latent space. Figure 3 reveals that the images generated by this procedure closely resemble those produced using the original sampling method. This confirms that tokens close in latent space can be treated as interchangeable, allowing for flexible token replacement without significantly compromising the visual semantics or overall image quality.
@@ 3.2 LANTERN: Relaxation of Acceptance Condition
We start with revisiting the original acceptance condition from Leviathan et al. (2023). The drafter model with distribution $p(x)$ samples a candidate token $\widetilde{x}\sim p(x|s)$ given a preceding sequence $s=(x_{1},\ldots,x_{n})$. The candidate is accepted with probability
$\min\left(1,\frac{q(\widetilde{x}|s)}{p(\widetilde{x}|s)}\right)$ (1)
where $q(x)$ is the target model’s distribution. If rejected, the next token is re-sampled from $[q(\cdot|s)-p(\cdot|s)]_{+}$ where $[\cdot]_{+}$ denotes normalization. Acceptance depends on the alignment of probabilities between the drafter and target models.
@@
Because $\widetilde{x}$ is always included in $B_{k}(\widetilde{x})$, this new accept probability is guaranteed to be equal to or higher than the original acceptance condition (1). As demonstrated in Table 2, applying LANTERN significantly increases the average acceptance probability, reaching values as high as 0.37. This improvement allows us to recover candidate tokens that would have otherwise been unjustifiably rejected.
@@ 3.3 With Limited Distributional Divergence
Table 2: Average accept probabilities of LANTERN. The accept probabilities are averaged on 100 image generations using text-conditioned LlamaGen Stage I model (Sun et al., 2024) and MS-COCO validation captions (Lin et al., 2014a). We only use accept probability of the first draft token. An average accept probability of EAGLE-2 (Li et al., 2024a) is 0.0402.
@@
For greedy decoding, LANTERN can be simply reduced to accept $\widetilde{x}$ if $\widetilde{x}=\operatorname*{arg\,max}_{x}q_{k,\delta}(x|s,D=\widetilde{x})$. Full algorithm can be found in Appendix B.
@@ 4 Experiments
In this section, we evaluate the performance of our method, LANTERN. In section 4.1, we report the experimental setup for our experiments. Section 4.2 evaluates LANTERN with other baselines in perspective of image quality and acceleration. In Section 4.3, we conduct an ablation study to clarify the effectiveness of each component in our method.
@@ 4.1 Experimental Setup
To validate our method LANTERN, we conduct experiments on LlamaGen-XL (Sun et al., 2024) for the text-conditional image generation as target model that demonstrate the best performance among autoregressive models without vision-specific modifications. We utilize the MS-COCO validation captions (Lin et al., 2014a) to generate images and evaluate the image quality with the ground-truth images. For the drafter model, we employ EAGLE-2 (Li et al., 2024a), which has demonstrated state-of-the-art performance in speculative decoding in the language domain. We employ the $\ell_{2}$ distance to quantify latent proximity, and utilize the TVD as a metric for divergence bound.
To assess the improvement of speed, we use $100$ MS-COCO validation captions to evaluate actual speedup and mean accepted length. Since measuring speedup with more than $100$ samples shows no significant difference, we use $100$ captions for efficiency. Actual speedup is measured by the inference time ratio between each method and vanilla auto-regressive decoding. Mean accepted length is determined by the average number of tokens accepted in each forward step of the target model. We evaluate each method in both the greedy decoding setting with $\tau=0$ and the sampling with $\tau=1$.
Since LANTERN can impact the quality of generated images, we evaluate image quality with FID (Heusel et al., 2017) and CLIP score (Hessel et al., 2021) with 30k samples. For each evaluation, Note that existing speculative decoding theoretically guarantees exact distribution matching with the target model, we do not evaluate image quality on EAGLE-2. Further details can be found in Appendix C.

### v3

@@ Abstract
Auto-Regressive (AR) models have recently gained prominence in image generation, often matching or even surpassing the performance of diffusion models. However, one major limitation of AR models is their sequential nature, which processes tokens one at a time, slowing down generation compared to models like GANs or diffusion-based methods that operate more efficiently. While speculative decoding has proven effective for accelerating LLMs by generating multiple tokens in a single forward, its application in visual AR models remains largely unexplored. In this work, we identify a challenge in this setting, which we term token selection ambiguity, wherein visual AR models frequently assign uniformly low probabilities to tokens, hampering the performance of speculative decoding. To overcome this challenge, we propose a relaxed acceptance condition referred to as LANTERN that leverages the interchangeability of tokens in latent space. This relaxation restores the effectiveness of speculative decoding in visual AR models by enabling more flexible use of candidate tokens that would otherwise be prematurely rejected. Furthermore, by incorporating a total variation distance bound, we ensure that these speed gains are achieved without significantly compromising image quality or semantic coherence. Experimental results demonstrate the efficacy of our method in providing a substantial speed-up over speculative decoding. In specific, compared to a naïve application of the state-of-the-art speculative decoding, LANTERN increases speed-ups by $\mathbf{1.75}\times$ and $\mathbf{1.82}\times$, as compared to greedy decoding and random sampling, respectively, when applied to LlamaGen, a contemporary visual AR model. The code is publicly available at https://github.com/jadohu/LANTERN.
@@ 1 Introduction
Auto-Regressive (AR) models have recently gained significant traction in image generation (Ramesh et al., 2021; Chen et al., 2020; Tian et al., 2024; Sun et al., 2024) due to their competitive performance, often matching or even surpassing diffusion models (Ho et al., 2020; Rombach et al., 2022). Notable examples include iGPT (Chen et al., 2020), DALL-E (Ramesh et al., 2021), VAR (Tian et al., 2024), and LlamaGen (Sun et al., 2024), which showcase the potential of AR models in image generation. Moreover, recent studies like Lu et al. (2023); Team (2024); Chern et al. (2024) have demonstrated that AR modeling can handle multi-modal data, including language and images, within a single unified framework. Given the remarkable success of AR models in language modeling, leading to the era of large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023), it is anticipated that AR modeling will emerge as a dominant paradigm for unifying multiple modalities into a single model in the near future.
@@
Figure 1: Images generated by vanilla decoding (top) and lossy speculative decoding with our relaxed acceptance condition (bottom) on the text-conditioned LlamaGen-XL Stage II (Sun et al., 2024). The mean accepted length for each image is displayed in white at the bottom right corner of each image.
@@
  • Our experiments using LlamaGen (Sun et al., 2024) as the target and EAGLE-2 (Li et al., 2024a) as the base speculative decoding method demonstrate significant speed-ups, improving from 1.29×1.29$\times$1.29 × to $\mathbf{2.26}\times$ in greedy decoding and from 0.93×0.93$\times$0.93 × to $\mathbf{1.69}\times$ in random sampling, compared to the naive application of EAGLE-2, without substantial performance drop in terms of image quality.
2 Preliminaries
Notations
In this paper, the target model refers to the visual AR model we aim to accelerate. In contrast, the drafter model is a supplementary model used to generate draft tokens. The probability distributions modeled by the drafter and target models are represented by $p(\cdot|\cdot)$ and $q(\cdot|\cdot)$, respectively. Individual tokens are denoted in lowercase $x$, and sequences are represented by uppercase $X$; for instance, $X_{i:j}$ represents the sequence $(x_{i},\ldots,x_{j})$. The concatenation of sequences $X_{1:N}$ and $Y_{1:M}$ is denoted by $(X_{1:N},Y_{1:M})=(x_{1},\ldots,x_{N},y_{1},\ldots,y_{M})$. For simplicity, we allow certain notational liberties; for instance, expressions like $X_{1:0}$ are treated as empty sequences to avoid unnecessary complexity in notation.
Visual Autoregressive Modeling
In visual AR models, image generation involves two main stages: generating image tokens through auto-regression and decoding the image tokens into actual image patches. In a text-to-image generation setting, given a tokenized text prompt $X_{1:N}$, the model generates a sequence of image tokens $X_{N+1:N+K}$ based on the following probability modeling:
$P(X_{N+1:N+K}\mid X_{1:N})=\prod_{\ell=1}^{K}P(x_{N+\ell}\mid X_{1:N+\ell-1}),$
where $K$ represents the total number of image tokens corresponding to the height and width of the image feature map. Since each token is predicted based solely on its preceding tokens, visual AR models require $K$ sequential steps to generate all $K$ image tokens.
Once the $K$ image tokens are generated, they are mapped to visual representation by referring to the codebook $\mathcal{C}$. Specifically, the codebook $\mathcal{C}=\{c_{1},\ldots,c_{L}\}$ consists of codes $c_{i}\in\mathbb{R}^{d}$, where each code is a latent feature used by the image encoder-decoder pair (such as VQVAE (Van Den Oord et al., 2017) or VQGAN (Esser et al., 2021)), and $d$ is the dimensionality of these features. As the text tokenization, since the image tokens represent indices in the codebook $\mathcal{C}$, each image token $x_{N+i}$ maps to $c_{x_{N+i}}$, which is then rearranged into a $h\times w\times d$ shaped tensor in raster-scan order (from top-left to bottom-right), for $h=H/f$ and $w=W/f$ when $f$ is down-sampling factor. After that, rearranged latent is fed into the image decoder $D:\mathbb{R}^{h\times w\times d}\rightarrow\mathbb{R}^{H\times W\times C}$ to construct an actual RGB image.
Speculative Decoding
We briefly introduce the basics of speculative decoding, which were proposed by Leviathan et al. (2023). Given a sequence of tokens $X_{1:N}=(x_{1},\ldots,x_{N})$, the drafter model generates $\gamma$ draft tokens $\widetilde{X}_{1:\gamma}$ as speculations for the next $\gamma$ tokens following $X_{1:N}$. Each draft token $\tilde{x}_{i}$ is sampled from the drafter distribution $p(x|(X_{1:N},\widetilde{X}_{1:i-1}))$ for each $i=1,\ldots,\gamma$, thus requiring $\gamma$ forward steps to generate $\gamma$ draft tokens.
After obtaining the draft tokens, the concatenated sequence $(X_{1:N},\widetilde{X}_{1:\gamma})$ is fed into the target model, which calculates the likelihood of each draft token $q(\tilde{x}_{i}|(X_{1:N},\widetilde{X}_{1:i-1})$ in parallel within a single forward pass. Each draft token $\tilde{x}_{i}$ is accepted with a probability:
$\min\left(1,\frac{q(\tilde{x}_{i}|(X_{1:N},\widetilde{X}_{1:i-1}))}{p(\tilde{x}_{i}|(X_{1:N},\widetilde{X}_{1:i-1}))}\right).$
If $\tilde{x}_{i}$ is accepted, it is immediately set as $x_{N+i}=\tilde{x}_{i}$. Otherwise, all subsequent tokens $\widetilde{X}_{i+1:\gamma}$ are discarded, and $x_{N+i}$ is resampled from a distribution defined by $[q(x|(X_{1:N},\widetilde{X}_{1:i-1}))-p(x|(X_{1:N},\widetilde{X}_{1:i-1}))]_{+}$, where $[\cdot]_{+}$ denotes normalization over positive values only. As proven by Leviathan et al. (2023), this approach ensures that the distribution of the generated token sequence matches the distribution produced by the target model. Related works on the visual AR models and speculative decoding are presented in Appendix A.
@@ 2 Preliminaries
3 Token Selection Ambiguity Limits Speculative Decoding in Visual AR Models
Although speculative decoding has been highly successful in LLMs, it fails to deliver comparable speed improvements when naïvely applied to visual AR models. As illustrated in Figure 2(a), the visual AR model (LlamaGen-3B (Sun et al., 2024)) exhibits a $41\%$ to $58\%$ reduction in mean accepted length compared to their performance in LLM (Vicuna-7B (Zheng et al., 2023)), a consistent decline observed across different speculative decoding methods. Toward an in-depth exploration of this performance degradation, we train drafters for various visual AR models and conduct additional analyses on their predictions.
Our analysis reveals that the drafter in visual AR models (LlamaGen-XL and Anole (Chern et al., 2024)) struggles to predict the target model’s outputs accurately. Specifically, as demonstrated in Figure 2(b), the trained drafters for visual AR models fail to capture the target model’s prediction precisely, whereas the drafter in LLMs exhibits considerably higher accuracy. Such a low accuracy of the drafters for visual AR models leads to a significant reduction in the speed-up achievable via speculative decoding.
To delve into the root cause of the drafter’s performance degradation, we analyze the next token probabilities of visual AR models and identify a unique problem we term token selection ambiguity. As shown in Figure 2(c), visual AR models present substantially lower average top-$1$ and top-$10$ probabilities in next token predictive distributions compared to language models, indicating that visual AR models are more ambiguous when selecting the next token. This lack of prioritization among tokens reflects the model’s limited confidence in any single option.
We hypothesize that this issue arises from fundamental differences between image and language data. In language models, tokens represent discrete units, such as words or subwords, that form structured and predictable sequences governed by grammar and syntax (Zipf, 1935). Consequently, the next-token probabilities are more concentrated, and the model has high confidence in the most likely token. In contrast, visual AR models treat pixels or patches as tokens, forming continuous and highly complex sequences. As a result, these models exhibit more dispersed next token probabilities and face higher uncertainty and ambiguity in predicting subsequent tokens. This token selection ambiguity problem in visual AR models hinders the drafter’s predictive accuracy, thereby limiting the effectiveness of speculative decoding. Further details on the setup of experiments can be found in Appendix B.1.
(a) Naïve Application Performance
(b) Drafter Test Accuracy
@@ 3 Token Selection Ambiguity Limits Speculative Decoding in Visual AR Models
Figure 2: (a) Mean accepted length of naïve application of existing speculative decoding methods on visual AR model and LLM counterpart. (b) Top-1 and top-3 accuracy of learned drafter model for predicting the target model’s outputs. (c) An average top-1 and top-10 probabilities in the next token prediction.
@@ 4.1 Latent Proximity Permits Token Interchangeability
This property arises from the tokenization process unique to visual AR models. Unlike text tokenization, which is straightforward due to its discrete nature, images are spatially continuous, making tokenization more complex Esser et al. (2021). To handle this, models like VQVAEs (Van Den Oord et al., 2017) and VQGANs (Esser et al., 2021) are used to discretize the latent embeddings of images, as introduced in Section 2. These embeddings maintain a continuous mapping between changes in latent space and the visual semantics of the generated images (Kingma & Welling, 2022; Goodfellow et al., 2014; Karras et al., 2019). As a result, small shifts in latent space lead to minor shifts in the image, supporting the idea that tokens close in the latent space are effectively interchangeable.
Figure 3: Image generated by text-conditioned LlamaGen-XL Stage II model (Sun et al., 2024). The images are generated by either standard sampling method (top) or sampling with random replacement within 100-closest tokens in the latent space (bottom).
To demonstrate this concept empirically, we perform an experiment in which, after each token is sampled, it is re-sampled uniformly from the 100 closest tokens in latent space. Figure 3 reveals that the images generated by this procedure closely resemble those produced using the original sampling method. This confirms that tokens close in latent space can be treated as interchangeable, allowing for flexible token replacement without significantly compromising the visual semantics or overall image quality. A more detailed analysis of latent proximity can be found in Appendix C.
@@ 4.2 LANTERN: Relaxation of Acceptance Condition
We start with revisiting the original acceptance condition from Leviathan et al. (2023), introduced in Section 2. The drafter model samples a draft token $\widetilde{x}\sim p(x|s)$ given a preceding sequence $s=X_{1:N}$. The draft token is accepted with probability $\min\left(1,\frac{q(\widetilde{x}|s)}{p(\widetilde{x}|s)}\right)$ and if rejected, the next token is re-sampled from $[q(x|s)-p(x|s)]_{+}$. Note that the acceptance depends on the alignment of probabilities between the drafter and target models.
@@
Because $\widetilde{x}$ is always included in $B_{k}(\widetilde{x})$, this new accept probability is guaranteed to be equal to or higher than the original acceptance condition. As demonstrated in Table 1, applying LANTERN significantly increases the average acceptance probability, reaching values as high as 0.37. This improvement allows us to recover candidate tokens that would have otherwise been unjustifiably rejected.
@@ 4.3 With Limited Distributional Divergence
Table 1: Average accept probabilities of LANTERN. We only use accept probability of the first draft token. An average accept probability of EAGLE-2 (Li et al., 2024a) is 0.0402.
@@
For greedy decoding, LANTERN can be simply reduced to accept $\widetilde{x}$ if $\widetilde{x}=\operatorname*{arg\,max}_{x}q_{k,\delta}(x|s,D=\widetilde{x})$. The algorithms for LANTERN and construction of $A_{k,\delta}$ can be found in Appendix D.
@@ 5.1 Experimental Setup
To validate our method LANTERN, we conduct experiments on text-conditioned LlamaGen-XL Stage I model (Sun et al., 2024) as a target model that establishes the best performance among visual AR models without vision-specific modifications. We utilize the MS-COCO validation captions (Lin et al., 2014) to generate images and evaluate the image quality with the ground-truth images. For the base speculative decoding method, we employ EAGLE-2 (Li et al., 2024a), which has demonstrated state-of-the-art performance in speculative decoding in the language domain. We employ the $\ell_{2}$ distance to quantify latent proximity and utilize the TVD as a metric for divergence bound.
For the assessment of speed-ups, we use $1000$ MS-COCO validation captions because this sample size provides results that are nearly identical to those obtained when evaluating the entire dataset. Actual speed-up is measured by the inference time ratio between each method and standard auto-regressive decoding. The mean accepted length is determined by the average number of tokens accepted in each forward step of the target model. We evaluate each method in both the greedy decoding setting with $\tau=0$ and the sampling with $\tau=1$. The statistical analysis on actual speed-up and number of captions can be found in Appendix F.1.
Since LANTERN can impact the quality of generated images, we evaluate image quality using FID (Heusel et al., 2017), CLIP score (Hessel et al., 2021), Precision and Recall (Kynkäänniemi et al., 2019), and HPS v2 (Wu et al., 2023) with 30K samples. For each evaluation, we do not evaluate image quality on EAGLE-2 since it theoretically guarantees exact distribution matching with the target model. Further details can be found in Appendix B.2.